A vector A has an x component Ax =3 and a y component Ay= -1. What are the x and y components of the vector 2A? The vector -4A?
1
Expert's answer
2020-10-02T07:14:51-0400

The vector can be written as:


"\vec{A} = (3,-1)"

Then


"2\vec{A} = 2\cdot (3,-1) = (2\cdot 3,2\cdot(-1)) = (6,-2)""-4\vec{A} = (-4)\cdot (3,-1) = ((-4)\cdot 3,(-4)\cdot(-1)) = (-12,4)"

Answer. "2\vec{A} = (6,-2)" and "-4\vec{A} =(-12,4)".